OpenStudy (anonymous):

What value of m solves the equation 2^m = 1/8

OpenStudy (irishboy123):

use \(log_2\)

OpenStudy (anonymous):

?

OpenStudy (irishboy123):

\(log_2 4 = 2\)

OpenStudy (anonymous):

To explain this with words, think of 2^m = 8. This would be 3, no? Then to get to 1/8, simply make the exponent negative.

OpenStudy (irishboy123):

\(log_a b=c \implies a^c = b\)

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

with what

OpenStudy (anonymous):

study island

OpenStudy (jhannybean):

\[2^m =\frac{1}{8}\]You need to solve for m, thus eliminating the base, taking \(\log_2\) of both sides, as @IrishBoy123 said would simplify the left side, therefore solving for m. \(\log_2(2) = 1\) \[\log_2(2^m) = \log_2\left(\frac{1}{8}\right)\]

OpenStudy (anonymous):

so the answer is 2

OpenStudy (anonymous):

you are you he,lp me

OpenStudy (anonymous):

hello

OpenStudy (freckles):

hint: \[\frac{1}{8}=\frac{1}{2^3}=2^{-3}\]